OpenStudy (anonymous):

Determine how much time is required for an investment to triple if interest is earned at the rate of 4.25% compounded weekly(52 weeks in a year)? given the formula y=a(1+r/n)^nt

OpenStudy (anonymous):

put \[3=(1+\frac{.0425}{52})^{52t}\] and solve for t

OpenStudy (anonymous):

really, 52 weeks in a year? thankfully they told us

OpenStudy (anonymous):

can you solve this or do you need steps?

OpenStudy (anonymous):

What satellite said (both parts).

OpenStudy (anonymous):

I think I have this one so no steps are needed thank you

OpenStudy (anonymous):

we go right to the answer you solve \[b^x=y\] by \[x=\frac{\ln(y)}{\ln(b)}\]

OpenStudy (anonymous):

44.85

OpenStudy (anonymous):

so \[52t=\frac{\ln(1+\frac{.0425}{52})}{\ln(3)}\] and \[t=\frac{\ln(1+\frac{.0425}{52})}{52\ln(3)}\] and